### v3

# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

Abstract
Accents are crucial in human communication as they help us understand others and allow us to communicate intelligibly in a way others understand us. While there has been significant progress in ASR, African-accented ASR has been understudied due to a lack of training datasets which are often expensive to create and demand colossal human labor. Our study aims to address this problem by automating the annotation process and reducing annotation-related expenses through informative uncertainty-based data selection. We propose a new multi-rounds adaptation process that uses epistemic uncertainty and evaluate it across several domains, datasets, and high-performing ASR models. Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. Moreover, the results of our active learning experiments, simulating real-world settings, where there are no gold transcriptions available, also demonstrate the ability of our approach to favor good quality real-life transcriptions. This indicates that our proposed approach addresses the immediate issue of African-accented ASR and has broader implications for improving ASR systems for other underrepresented and low-resource languages and accents. We open-source the code here.
Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
Bonaventure F. P. Dossou1,2,3,4∗, Atnafu Lambebo Tonja5,∗, Chris Chinenye Emezue1,4,6,∗, Tobi Olatunji7,8,∗ Naome A Etori9,∗, Salomey Osei10,∗, Tosin Adewumi11,∗, Sahib Singh12,∗ ∗Masakhane NLP, 1Mila Quebec AI Institute, 2 McGill University, 3 Lelapa AI, 4 Lanfrica 5Instituto Politécnico Nacional, 6Technical University of Munich, 7Intron Health Inc, 8Georgia Institute of Technology 9University of Minnesota - Twin Cities, 10Universidad de Deusto, 11 Luleå University of Technology, 12 Ford Motor Company
1 Introduction
Generally speaking, Accented English is a form of spoken English that exhibits intonations and pronunciations different from standard English. For instance, Nigerian English differs from British English, which differs from American English. In this work, we consider African-accented English as a form of English spoken by African-native speakers.
Automatic Speech Recognition (ASR) is an active area of research Kodish-Wachs et al. (2018); Finley et al. (2018); Zapata and Kirkedal (2015). In the African context, rich in linguistic diversity and complexity, ASR systems have great potential to reduce language barriers in daily communication. In particular, there is a considerable interest in the development of ASR systems that can accurately distinguish African-accented speech, with a focus on creating systems that can accurately transcribe and interpret speech in a variety of accents and dialects (Koenecke et al., 2020; Hirayama et al., 2012; Tadimeti et al., 2022; Ali et al., 2015; Besacier et al., 2014; Yemmene and Besacier, 2019; Gauthier et al., 2016; Reitmaier et al., 2022; Afonja et al., 2021b). However, developing and deploying ASR systems in an African setting is a challenging task Yemmene and Besacier (2019); De Vries et al. (2014) due to the diversity of African languages, with several dialects and variants characterized by low-resource environments and limited data availability Badenhorst and De Wet (2019, 2017); Barnard et al. (2009); Yemmene and Besacier (2019); DiChristofano et al. (2022). Our methodology facilitates the development of robust, generalizable, and cost-efficient African-accented ASR models. The enhanced generalization for out-of-distribution (OOD) cases, particularly for accents with minimal resources, suggests that our approach extends beyond addressing the specific challenges of African-accented ASR. It holds promise for enhancing ASR systems for other marginalized and resource-scarce languages and accents. This advancement could be a crucial step towards establishing more inclusive and globally accessible ASR technologies. In this work, we reduce this gap, through several contributions listed as follows:
using epistemic uncertainty, we propose a novel multi-rounds adaptation process of several high-performing pretrained ASR models,
we evaluate our approach across several speech domains (clinical, general, both), African-accented speech datasets: AfriSpeech-200 Olatunji et al. (2023), SautiDB Afonja et al. (2021b), MedicalSpeech and CommonVoices English Accented Dataset Ardila et al. (2019), and provide domain-specific as well as accent-specific analyses,
we show that our approach improves the WER performance by 69.44% while reducing significantly the required amount of labeled data (by 35-45%),
we show, based on additional active learning experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
2 Challenges and Related Works for African-accented ASR
Although the 2000+ African Languages make up about 31% of the 7000+ languages spoken in the world Eberhard et al. (2019), almost all African languages are classified as "low-resource" Tsvetkov (2017). Despite advancements in ASR over the recent years Kodish-Wachs et al. (2018); Blackley et al. (2019), research or applications in the African context are not well investigated in the literature, likely a product of the paucity of data for accented languages, inadequate research funding, data-hungry approaches well-suited to high-resource settings, and most importantly, data-efficient approaches are inadequately explored. As a result of scarce data and resources, it is also challenging to pretrain large multilingual language models for African languages Dossou et al. (2022).
Despite its potential to improve language communication, ASR models suffer from racial bias and poor performance on accented speech Koenecke et al. (2020), particularly African accents. Recognizing heavy accents in speech data is a challenge for ASR systems since it contains high variability in terms of pronunciations Das et al. (2021). In recent years, we have seen a drastic improvement in ASR systems due to more advanced deep learning systems and the availability of data. It is however shocking that racial disparities and poor performances in accented speech still exist for black speakers. Typically, ASR system generalization is hard to achieve due to difficulties in getting good speech transcript data, including for African accents. The different dialects and accents of many languages introduce some level of difficulty in transcribing such data. These variations are much more complicated in the case of African languages, as speech data in general is more difficult to source. For this reason, there is a poor performance of ASR in African languages.
Improving ASR on African accented speech and linguistic diversity is hence necessary to avoid speakers having to change their way of speaking Mengesha et al. (2021) (a situation that will likely make speakers feel excluded, and the social impact is high), while bringing enormous potential to transform several working areas and daily human needs and communication.
3 Datasets and Methodology
In this work, we established several baselines, across several high-performing ASR systems. Our baselines are produced by feeding the pretrained ASR models, the entirety (100%) of available labeled data. To investigate the effectiveness of our approach, we impose data constraints, choosing fewer training examples than available, to not exceed 60-65% of the entire initial dataset. This simulates a more realistic scenario where the entirety of the data might not be available or feasible to use, thus testing the robustness and efficiency of our method under constrained conditions. We utilize 30% of the initial dataset as our training set and set the remaining 70% as the pool dataset. The purpose of the pool dataset in our study is to allow the model, in our proposed framework, to learn to carefully select the most informative data samples.
To further solidify and demonstrate the effectiveness of our method, we leveraged it in an active learning (AL) setup. AL is a semi-supervised machine learning algorithm that makes use of only a few initial training data points to achieve better performance of a given model. It is important to note that due to the lack of African-accented ASR benchmarks, we primarily compared the results, to our baselines and our best-performing selection technique. Within the defined constraints, for our experiments, for our proposed method, we adjust the sizes of the training and pool datasets, based on the number of training examples available for each domain (clinical, general, and both) (see Tables 2, 4 and Appendix A).
3.1 Datasets
In this work, we primarily used the AfriSpeech-200 dataset Olatunji et al. (2023), a novel 200-hour African-accented English speech corpus carefully curated for clinical and general domains ASR. The dataset covers 120+ African accents, all derived from 5 language families Eberhard et al. (2019): Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across western, eastern, and southern Africa. AfriSpeech-200 is a crowd-sourced dataset involving over 2000 African speakers, representing 13 anglophone countries across sub-Saharan Africa and the US (see Table 1).
To establish the dataset-agnostic nature of our approach, we explore three additional datasets:
1.
SautiDB Afonja et al. (2021a): a dataset of Nigerian accent recordings with 919 audio samples, a sampling rate of 48kHz each, for a total amount of 59 minutes of recordings,
2.
MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent: a dataset which contains 6,661 audio utterances for common medical symptoms like knee pain or headache, for a total of 8 hours of recordings;
3.
CommonVoices English Accented Dataset: a subset of English CommonVoices (version 10) Ardila et al. (2019), with western accents removed as our primary interest is in low-resource settings.
| AfriSpeech Dataset Statistics | |
| Total duration | 200.91 hrs |
| Total clips | 67,577 |
| Unique Speakers | 2,463 |
| Average Audio duration | 10.7 seconds |
| Speaker Gender Ratios - # Clip % | |
| Female | 57.11% |
| Male | 42.41% |
| Other/Unknown | 0.48% |
| Speaker Age Groups - # Clips | |
| <18yrs | 1,264 (1.88%) |
| 19-25 | 36,728 (54.58%) |
| 26-40 | 18,366 (27.29%) |
| 41-55 | 10,374 (15.42%) |
| >56yrs | 563 (0.84%) |
| Clip Domain - # Clips | |
| Clinical | 41,765 (61.80%) |
| General | 25,812 (38.20%) |
| AfriSpeech-200 Dataset Splits | ||||
| Item | Train (*AL Train) | Dev | Test | AL Top-k |
| # Speakers | 1466 | 247 | 750 | ✗ |
| # Hours | 173.4 | 8.74 | 18.77 | ✗ |
| # Accents | 71 | 45 | 108 | ✗ |
| Avg secs/speaker | 425.81 | 127.32 | 90.08 | ✗ |
| clips/speaker | 39.56 | 13.08 | 8.46 | ✗ |
| speakers/accent | 20.65 | 5.49 | 6.94 | ✗ |
| secs/accent | 8791.96 | 698.82 | 625.55 | ✗ |
| # general domain | 21682 (*6504) | 1407 | 2723 | 2000 |
| # clinical domain | 36318 (*10895) | 1824 | 3623 | 3500 |
| # both domain | 58000 (*17400) | 3221 | 6346 | 6500 |
3.2 Methodology
Our methodology is based on epistemic uncertainty. Epistemic Uncertainty (EU) is the uncertainty of the model, most often due to a lack of training data. For a given input , EU is usually defined with the variance (for a given model with some parameters ) as follows:
where . is the learned approximation of . In our experiments, we decided to use the standard deviation defined (STD) as . EU is useful for the model’s training as it improves its robustness and encourages exploration and diversity, to mitigate inductive bias related to underrepresented accents (accents with fewer speech data and speakers).
One of the most popular ways to estimate uncertainty is by inferring predictive distributions with Bayesian Neural Networks (BNNs). A predictive distribution is defined as where is the target, is the input and is the dataset. One can thus inspect the variance (or the standard deviation, as we used in our experiments) and uncover uncertainty, using the predictive distribution. A way of learning the predictive distribution is to learn a distribution over the parameters i.e. the parametric posterior distribution .
There are many ways of computing epistemic uncertainty, but the two most common approaches are Monte Carlo Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017). Many empirical results demonstrated that Deep Ensembles generally work better while being very computationally expensive. Due to computational resource constraints, we decided to demonstrate the effectiveness of our approach using Mc-Dropout empirically. Mc-Dropout provides a scalable way to learn a predictive distribution. It randomly switches off neurons acting as regularizers in a given neural network. Each dropout configuration corresponds to a different sample from the approximate parametric posterior distribution . Sampling from the (approximate) posterior enables Monte Carlo integration of the likelihood of the model, which uncovers the predictive distribution. Our uncertainty-based selection pipeline is detailed in Figure 1 and the algorithm 1. In each adaptation round, we use a finetuned model, to select the most uncertain samples from the pool dataset and add them to the training set for the next round of the model adaptation.
While performing AL experiments, we consider (mask) the samples from the pool dataset to be entirely unlabeled (i.e. no gold transcriptions as targets). This implies that we simulate an unsupervised context, and learn to select the best possible predicted transcription as the target (gold) transcriptions.
Using Mc-Dropout sets a pre-requisite for the model to be equipped with Dropout components during pretraining. This condition eliminates some high-level ASR models like Whisper whisper. However, we finetuned Whisper in our baseline setting and evaluated it against our best-performing approach. This is to assess its robustness and adequacy. In this work, we use the following four state-of-the-art (SOTA) ASR models: Wav2Vec2-XLSR-53 Conneau et al. (2020), Hubert-Large Hsu et al. (2021), WavLM-Large Chen et al. (2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al. (2020), which we hereafter refer to as Wav2Vec, Hubert, WavLM, and Nemo respectively.
3.3 Experimental Design
Added to the baselines described above, in our framework, we explored two selection techniques: most and random. The most technique consists of selecting the most uncertain audio samples from the pool dataset, while the random one, consists of selecting random audio samples across the pool dataset. In this work, we used the words "sampling", "techniques", and "selection" interchangeably with the same meaning.
We performed various experiments to evaluate our proposed approach’s efficiency and applicability. Each experiment at the model level is performed in three domains: general, clinical, and both (which is the combination of general and clinical data).
For each domain, we have four setups: (a) baseline which consists of simply finetuning a pretrained model on the entire dataset (as described in section 3.1), (b) EU-Most where the pretrained model is finetuned within our framework using the most selection, (c) Random uses the random selection, and (d) AL (w/ EU-Most) where the pretrained model is fine-tuned using active learning and the most technique.
Our experiments are designed to allow us to answer the following specific questions:
1.
given a set of African accents , how does the pretrained ASR model adapt to across the adaptation rounds and domains (clinical, general and both)?
2.
which selection strategy between most and random works better, and for which domain(s)?
3.
which domain(s) help(s) the model to do better, and how does the model perform (in terms of uncertainty) across the domain(s)?
4.
is epistemic uncertainty-based selection effective in low-resource data scenarios?
5.
is uncertainty-based selection, model, and dataset agnostic?
To answer questions 1-4, we define a metric called Uncertainty WER (U-WER). U-WER is described as follows:
where is the speech input, the gold target (transcript), the finetuned ASR model. Due to the nature of the MC-Dropout technique, running it times will produce different speech representations for the original speech input . Each speech representation is denoted , which could consequently lead to different predicted transcription, that we denote by . is the WER function, computing the WER between the predicted transcript and the gold (true) transcript . is the average of WERs between the different speech representations and the target prediction. In our experiments, . The ideal case would be when U-WER0, because this would mean that the model is robust enough to produce consistent transcripts and speech representation instances .
The rationale behind the U-WER is the following: as we acquire new data points that are intrinsically beneficial to the model and its learning, the U-WER should (consistently) decrease or be constant across the multiple adaptation rounds. We could therefore affirm that the model is gaining in robustness, knowledge, and performance; which is a very important aspect of generalization. This last point will be enforced by answering question 5 for which we used the standard WER.
To answer question 5, we evaluated our approach with three additional ASR pretrained models (Nemo, WavLM, and Hubert), and across 3 external datasets (SautiDB, Common Voice English Accented, and MedicalSpeech). Additionally, to answer questions 1-4, for consistency and better visualization, we consider exclusively the top-10 (in terms of frequency) accents present across three adaptation rounds and both selection methods. Regarding the very low-resource settings, we consider the 5 accents with the least number of hours of recordings.
In our AL setup, the U-WER is computed using pairwise WER scores among the predicted transcriptions, rather than using the gold transcriptions. To select the ideal or best predicted transcript for the unlabeled speech , we proceed as described in Algorithm 2.
This implies that for the current speech input , is computed by taking the standard deviation ‘wer_list‘ variable (this corresponds exactly to the formula of the described above). Finally, we assign to , the predicted transcription with the lowest average WER i.e. the lowest value of the dictionary ‘wer_target_dict‘. For our experiments, we use various types of GPUs: 6 RTX 800 (with 48 and 128 GB), and 4 A100 (with 48 and 128 GB). Our training and evaluation were constantly deployed over a month. Our models have 311281844 trainable parameters. Each audio sample has been normalized and processed with a sample rate of 16kHz.
| Model | General | Clinical | Both | |||||||||
| Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | |
| Wav2vec | 0.4980 | 0.1111 | 0.1011 | 0.1059 | 0.5610 | 0.3571 | 0.2457 | 0.2545 | 0.5300 | 0.1666 | 0.1266 | 0.1309 |
| **Hubert | 0.1743 | ✗ | 0.1901 | 0.1887 | 0.2907 | ✗ | 0.2594 | 0.2709 | 0.2365 | ✗ | 0.2453 | 0.2586 |
| **WavLM | 0.1635 | ✗ | 0.1576 | 0.1764 | 0.3076 | ✗ | 0.2313 | 0.2537 | 0.2047 | ✗ | 0.1897 | 0.1976 |
| *Nemo | 0.2824 | ✗ | 0.1765 | 0.1815 | 0.2600 | ✗ | 0.2492 | 0.2526 | 0.3765 | ✗ | 0.2576 | 0.2610 |
| Average Performance | 0.2796 | - | 0.1563 | 0.1631 | 0.3548 | - | 0.2464 | 0.2579 | 0.3369 | - | 0.2043 | 0.2120 |
| Whisper-Medium Olatunji et al. (2023) | 0.2806 | - | - | - | 0.3443 | - | - | - | 0.3116 | - | - | - |
| Dataset | Split and Size for our approach | Finetuning Epochs | Baseline | Uncertainty Sampling w/ most | |||
| Train | Aug | Top-k | Test | (Entire training dataset) | (Train + Aug) | ||
| SautiDB Afonja et al. (2021a) | 234 | 547 | 92 | 138 | 50 | 0.50 | 0.12 |
| MedicalSpeech | 1598 | 3730 | 1333 | 622 | 5 | 0.30 | 0.28 |
| CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) | 26614 | 62100 | 10350 | 232 | 5 | 0.50 | 0.22 |
| Average | ✗ | ✗ | ✗ | ✗ | ✗ | 0.43 | 0.20 |
4 Results and Discussion
To compute the performance improvement for each domain, we take the ratio of the average of our best method and the baseline. Table 3 presents the results of our experiments. We observe that our uncertainty-based selection approach yields considerable performance gains over the baselines, across all models, all domains, and datasets: general (55.90%), clinical (69.44), and both (61.74%). Additionally, our approach significantly outperforms Whisper-Medium; affirming the effectiveness and reliability of our method. This shows the importance and usefulness of epistemic uncertainty in ASR for low-resource languages.
More importantly, we observe that the most selection provides the best results across all domains. The predominance of most makes sense because, through continual and progressive exposure to the most uncertain samples, the model updates its parameters to become more robust to the noise from uncertain samples. Therefore, it learns to represent those uncertain samples better, yielding a performance improvement. Nevertheless, we speculate that gaps in performance between the general and clinical domains are because clinical samples are the most difficult to capture due to various factors (e.g., domain-specific clinical jargon). This result and analysis provide an answer to question 2, establishing most technique as the most beneficial selection strategy across all domains (see also Figures 2, 3 and 4).
Given the diversity of traits (accents, speakers, gender, age) across the dataset, we attempt to characterize the audio samples and domains that provide the best learning signal to the model. To ground our analysis, we evaluated accent information for top-k uncertain samples. The results are presented in Figures 2, 3 and 4. For further analyses, we focus only on the results of the most technique. For each domain, we find that the top-10 accents are consistent across all rounds (see Figures 2 (a), 3 (a), and 4 (a)), representing the accents with the top-10 highest number of recording hours. With each successive most uncertain top-k samples, we observe that those accents remain the most uncertain and therefore the most difficult to transcribe. From the country and accent statistics (Tables 6, 7 and 8), we observe that those languages have the highest number of speakers and geographical spread across many countries. This interesting observation can be understood as the linguistic richness factor, which potentially explains the high variance induced in the dataset. This implies that, in general, the accents that benefit the model and its learning are the ones that are linguistically rich and diverse. Generally speaking, the model adapts and continuously improves its performance on those accents across the rounds. It is important to notice that for the clinical and both domains (see Figures 3 (a), and 4 (a)), the accent level analysis also reveals some perturbation and noise in the WER variation, which remains constant or slightly increases from Round 2 to Round 3: this suggests that for a critical and complex domain (like the clinical one), there is, and should be, a trade-off between linguistic diversity, rounds, and the value of k samples to be selected. Ultimately, we can positively answer questions 1 and 3, affirming that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also provide a positive answer to question 4, as we can see a generally improving (or steady) performance on truly low-resource accents in our training dataset, across domains and rounds. This makes our approach very relevant to the low-resourcefulness situation prevalent in many African accents and languages.
Finally, to demonstrate the agnostic aspect of our approach, we evaluated it with three different pretrained ASR models (Hubert, WavLM, and Nemo), and with three datasets containing accented speech on general and clinical domains, only using most selection method. The results are respectively presented in Tables 3 and 4. The main takeaway is that our uncertainty-based adaptation approach performs better than baselines. This efficiently demonstrates that our approach can be applied to any model architecture and with any dataset.
5 Conclusion
We introduced an epistemic uncertainty-based adaptation process for developing efficient, robust African-accented ASR models that perform well across various models, domains, and datasets, particularly in low-resource settings. Our approach effectively identifies and leverages the most informative data samples, enhancing model robustness across different domains.
6 Limitations
6.1 Trade-off between adaptation rounds and new data points selected
In section 4, the results and our analyses revealed that while our approach provided performance improvements, and benefited from linguistically rich accents, we observed that for complex domains (such as clinical and both), it is important to have a stopping criteria and a trade-off between the number of adaptation rounds and the amount of new data points selected at each round. We believe doing such analyses will provide useful insight to improve our current approach.
6.2 Computational Complexity
Similar to machine learning and deep learning approaches, this work was limited by the availability of and access to computational resources. With more computational resources, we would explore Deep Ensembles Lakshminarayanan et al. (2017) as an alternative to our current approach (Mc-Dropout) to estimate epistemic uncertainty from the model.
Appendix A Hyper-parameters
Table 5 shows the hyper-parameter settings used in this study. The top-k value in the table is changed according to the domain used in each of the experiments. For example, when conducting experiments in the general domain, we set the value of top-k to 2k.
| Hyper-parameters | Values |
| attention dropout | 0.1 |
| hidden dropout | 0.1 |
| layer drop | 0.1 |
| train batch size | 16 |
| val batch size | 8 |
| number of epochs | 5 |
| learning rate | 3e-4 |
| maximum audio length | 260000 |
| maximum label length | 260 |
| minimum transcript length | 10 |
| top_k | 2000, 3500, 6500 |
| domains | general, clinical, all |
| active learning rounds | 3 |
| sampling mode | most, random |
| mc-dropout round | 10 |
Appendix B Country Statistics
Table 6 shows the statistics of the countries across the AfriSpeech-200 dataset.
| Country | Clips | Speakers | Duration (seconds) | Duration (hrs) |
| Nigeria | 45875 | 1979 | 512646.88 | 142.40 |
| Kenya | 8304 | 137 | 75195.43 | 20.89 |
| South Africa | 7870 | 223 | 81688.11 | 22.69 |
| Ghana | 2018 | 37 | 18581.13 | 5.16 |
| Botswana | 1391 | 38 | 14249.01 | 3.96 |
| Uganda | 1092 | 26 | 10420.42 | 2.89 |
| Rwanda | 469 | 9 | 5300.99 | 1.47 |
| United States of America | 219 | 5 | 1900.98 | 0.53 |
| Turkey | 66 | 1 | 664.01 | 0.18 |
| Zimbabwe | 63 | 3 | 635.11 | 0.18 |
| Malawi | 60 | 1 | 554.61 | 0.15 |
| Tanzania | 51 | 2 | 645.51 | 0.18 |
| Lesotho | 7 | 1 | 78.40 | 0.02 |
Appendix C Dataset Accents Stats
Tables 7 and 8 provide a list of AfriSpeech accents along with the number of unique speakers, countries where speakers for each accent are located, duration in seconds for each accent, and their presence in the train, dev, and test splits.
| Accent | Clips | Speakers | Duration(s) | Countries | Splits |
| yoruba | 15407 | 683 | 161587.55 | US,NG | train,test,dev |
| igbo | 8677 | 374 | 93035.79 | US,NG,ZA | train,test,dev |
| swahili | 6320 | 119 | 55932.82 | KE,TZ,ZA,UG | train,test,dev |
| hausa | 5765 | 248 | 70878.67 | NG | train,test,dev |
| ijaw | 2499 | 105 | 33178.9 | NG | train,test,dev |
| afrikaans | 2048 | 33 | 20586.49 | ZA | train,test,dev |
| idoma | 1877 | 72 | 20463.6 | NG | train,test,dev |
| zulu | 1794 | 52 | 18216.97 | ZA,TR,LS | dev,train,test |
| setswana | 1588 | 39 | 16553.22 | BW,ZA | dev,test,train |
| twi | 1566 | 22 | 14340.12 | GH | test,train,dev |
| isizulu | 1048 | 48 | 10376.09 | ZA | test,train,dev |
| igala | 919 | 31 | 9854.72 | NG | train,test |
| izon | 838 | 47 | 9602.53 | NG | train,dev,test |
| kiswahili | 827 | 6 | 8988.26 | KE | train,test |
| ebira | 757 | 42 | 7752.94 | NG | train,test,dev |
| luganda | 722 | 22 | 6768.19 | UG,BW,KE | test,dev,train |
| urhobo | 646 | 32 | 6685.12 | NG | train,dev,test |
| nembe | 578 | 16 | 6644.72 | NG | train,test,dev |
| ibibio | 570 | 39 | 6489.29 | NG | train,test,dev |
| pidgin | 514 | 20 | 5871.57 | NG | test,train,dev |
| luhya | 508 | 4 | 4497.02 | KE | train,test |
| kinyarwanda | 469 | 9 | 5300.99 | RW | train,test,dev |
| xhosa | 392 | 12 | 4604.84 | ZA | train,dev,test |
| tswana | 387 | 18 | 4148.58 | ZA,BW | train,test,dev |
| esan | 380 | 13 | 4162.63 | NG | train,test,dev |
| alago | 363 | 8 | 3902.09 | NG | train,test |
| tshivenda | 353 | 5 | 3264.77 | ZA | test,train |
| fulani | 312 | 18 | 5084.32 | NG | test,train |
| isoko | 298 | 16 | 4236.88 | NG | train,test,dev |
| akan (fante) | 295 | 9 | 2848.54 | GH | train,dev,test |
| ikwere | 293 | 14 | 3480.43 | NG | test,train,dev |
| sepedi | 275 | 10 | 2751.68 | ZA | dev,test,train |
| efik | 269 | 11 | 2559.32 | NG | test,train,dev |
| edo | 237 | 12 | 1842.32 | NG | train,test,dev |
| luo | 234 | 4 | 2052.25 | UG,KE | test,train,dev |
| kikuyu | 229 | 4 | 1949.62 | KE | train,test,dev |
| bekwarra | 218 | 3 | 2000.46 | NG | train,test |
| isixhosa | 210 | 9 | 2100.28 | ZA | train,dev,test |
| hausa/fulani | 202 | 3 | 2213.53 | NG | test,train |
| epie | 202 | 6 | 2320.21 | NG | train,test |
| isindebele | 198 | 2 | 1759.49 | ZA | train,test |
| venda and xitsonga | 188 | 2 | 2603.75 | ZA | train,test |
| sotho | 182 | 4 | 2082.21 | ZA | dev,test,train |
| akan | 157 | 6 | 1392.47 | GH | test,train |
| nupe | 156 | 9 | 1608.24 | NG | dev,train,test |
| anaang | 153 | 8 | 1532.56 | NG | test,dev |
| english | 151 | 11 | 2445.98 | NG | dev,test |
| afemai | 142 | 2 | 1877.04 | NG | train,test |
| shona | 138 | 8 | 1419.98 | ZA,ZW | test,train,dev |
| eggon | 137 | 5 | 1833.77 | NG | test |
| luganda and kiswahili | 134 | 1 | 1356.93 | UG | train |
| ukwuani | 133 | 7 | 1269.02 | NG | test |
| sesotho | 132 | 10 | 1397.16 | ZA | train,dev,test |
| benin | 124 | 4 | 1457.48 | NG | train,test |
| kagoma | 123 | 1 | 1781.04 | NG | train |
| nasarawa eggon | 120 | 1 | 1039.99 | NG | train |
| tiv | 120 | 14 | 1084.52 | NG | train,test,dev |
| south african english | 119 | 2 | 1643.82 | ZA | train,test |
| borana | 112 | 1 | 1090.71 | KE | train |
| Accent | Clips | Speakers | Duration(s) | Countries | Splits |
| swahili ,luganda ,arabic | 109 | 1 | 929.46 | UG | train |
| ogoni | 109 | 4 | 1629.7 | NG | train,test |
| mada | 109 | 2 | 1786.26 | NG | test |
| bette | 106 | 4 | 930.16 | NG | train,test |
| berom | 105 | 4 | 1272.99 | NG | dev,test |
| bini | 104 | 4 | 1499.75 | NG | test |
| ngas | 102 | 3 | 1234.16 | NG | train,test |
| etsako | 101 | 4 | 1074.53 | NG | train,test |
| okrika | 100 | 3 | 1887.47 | NG | train,test |
| venda | 99 | 2 | 938.14 | ZA | train,test |
| siswati | 96 | 5 | 1367.45 | ZA | dev,train,test |
| damara | 92 | 1 | 674.43 | NG | train |
| yoruba, hausa | 89 | 5 | 928.98 | NG | test |
| southern sotho | 89 | 1 | 889.73 | ZA | train |
| kanuri | 86 | 7 | 1936.78 | NG | test,dev |
| itsekiri | 82 | 3 | 778.47 | NG | test,dev |
| ekpeye | 80 | 2 | 922.88 | NG | test |
| mwaghavul | 78 | 2 | 738.02 | NG | test |
| bajju | 72 | 2 | 758.16 | NG | test |
| luo, swahili | 71 | 1 | 616.57 | KE | train |
| dholuo | 70 | 1 | 669.07 | KE | train |
| ekene | 68 | 1 | 839.31 | NG | test |
| jaba | 65 | 2 | 540.66 | NG | test |
| ika | 65 | 4 | 576.56 | NG | test,dev |
| angas | 65 | 1 | 589.99 | NG | test |
| ateso | 63 | 1 | 624.28 | UG | train |
| brass | 62 | 2 | 900.04 | NG | test |
| ikulu | 61 | 1 | 313.2 | NG | test |
| eleme | 60 | 2 | 1207.92 | NG | test |
| chichewa | 60 | 1 | 554.61 | MW | train |
| oklo | 58 | 1 | 871.37 | NG | test |
| meru | 58 | 2 | 865.07 | KE | train,test |
| agatu | 55 | 1 | 369.11 | NG | test |
| okirika | 54 | 1 | 792.65 | NG | test |
| igarra | 54 | 1 | 562.12 | NG | test |
| ijaw(nembe) | 54 | 2 | 537.56 | NG | test |
| khana | 51 | 2 | 497.42 | NG | test |
| ogbia | 51 | 4 | 461.15 | NG | test,dev |
| gbagyi | 51 | 4 | 693.43 | NG | test |
| portuguese | 50 | 1 | 525.02 | ZA | train |
| delta | 49 | 2 | 425.76 | NG | test |
| bassa | 49 | 1 | 646.13 | NG | test |
| etche | 49 | 1 | 637.48 | NG | test |
| kubi | 46 | 1 | 495.21 | NG | test |
| jukun | 44 | 2 | 362.12 | NG | test |
| igbo and yoruba | 43 | 2 | 466.98 | NG | test |
| urobo | 43 | 3 | 573.14 | NG | test |
| kalabari | 42 | 5 | 305.49 | NG | test |
| ibani | 42 | 1 | 322.34 | NG | test |
| obolo | 37 | 1 | 204.79 | NG | test |
| idah | 34 | 1 | 533.5 | NG | test |
| bassa-nge/nupe | 31 | 3 | 267.42 | NG | test,dev |
| yala mbembe | 29 | 1 | 237.27 | NG | test |
| eket | 28 | 1 | 238.85 | NG | test |
| afo | 26 | 1 | 171.15 | NG | test |
| ebiobo | 25 | 1 | 226.27 | NG | test |
| nyandang | 25 | 1 | 230.41 | NG | test |
| ishan | 23 | 1 | 194.12 | NG | test |
| bagi | 20 | 1 | 284.54 | NG | test |
| estako | 20 | 1 | 480.78 | NG | test |
| gerawa | 13 | 1 | 342.15 | NG | test |
Appendix D Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with both random and most selection methods.
Appendix E Ascending and Descending Accents
Figure 7 shows ascending and descending accents across the Top 2k most uncertain samples.