### v7

# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. By combining several active learning paradigms and the core-set approach, we propose a new multi-round adaptation process that utilizes epistemic uncertainty to automate annotation, thereby significantly reducing associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples that contribute most to model uncertainty, thereby enhancing training efficiency. We define a new U-WER metric to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 27% WER relative average improvement while requiring, on average, 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
Bonaventure F. P. Dossou McGill University, Mila Quebec AI Institute bonaventure.dossou@mila.quebec
1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication Kodish-Wachs et al. (2018); Finley et al. (2018); Zapata and Kirkedal (2015). Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents (Eberhard et al. (2019); Tsvetkov (2017)). This gap highlights the need for ASR systems that can effectively handle the linguistic diversity and complexity of African languages, particularly in critical applications such as healthcare. Due to the lack of representations of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names (Olatunji et al. (2023a)).
To address these challenges, our work focuses on adapting pre-trained speech models to transcribe African-accented English more accurately, characterized by unique intonations and pronunciations Benzeghiba et al. (2007); Hinsvark et al. (2021). We use epistemic uncertainty (EU) Kendall and Gal (2017) to guide the adaptation process by identifying gaps in model knowledge and prioritizing data for the model to learn from next. This is particularly beneficial in scenarios where data annotation is costly or time-consuming, as often seen in the African context Badenhorst and De Wet (2019, 2017); Barnard et al. (2009); Yemmene and Besacier (2019); DiChristofano et al. (2022); Dossou et al. (2022); Dossou and Emezue (2021). EU also improves robustness and encourages exploration to mitigate inductive bias from underrepresented accents. Common approaches to compute EU include Monte Carlo Dropout (MC-Dropout) Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017), with the latter being more effective but computationally expensive. Due to resource constraints, we utilize MC-Dropout, which necessitates that models incorporate dropout components during pretraining.
We employ Active Learning (AL) techniques further to enhance the efficiency and effectiveness of model adaptation. AL leverages epistemic uncertainty to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) Gal et al. (2017); Houlsby et al. (2011) and Adversarial Active Learning (AAL) Ducoffe and Precioso (2018). AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with complex cases to enhance robustness. The core-set approach (CSA) Sener and Savarese (2017) is also related, as it selects a subset of the training data to ensure that a model trained on this subset performs comparably to one trained on the entire dataset, thereby addressing scalability and efficiency. A critical component of AL is the acquisition function (AF), which determines the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) Liu and Li (2023), Bayesian Active Learning by Disagreement (BALD) Gal et al. (2017), and BatchBALD Kirsch et al. (2019). US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD that selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, making it ideal for efficient data labeling.
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order): First, we integrate the CSA by leveraging smaller training subsets ($\sim 45\%$ smaller than the entire available training sets). Second, we utilize DBAL with MC-Dropout to apply dropout during both training and inference, thereby estimating the Bayesian posterior distribution. This allows us to practically and efficiently estimate EU in the models used Gal et al. (2017) (see section 3.2 for more details). Third, we use the estimated EU and integrate the idea of AAL using the US acquisition function.
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 Olatunji et al. (2023b)), SautiDB Afonja et al. (2021b), MedicalSpeech, CommonVoices English Accented Dataset Ardila et al. (2019), and several high-performing speech models (Wav2Vec2-XLSR-53 Conneau et al. (2020), HuBERT-Large Hsu et al. (2021), WavLM-Large Chen et al. (2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al. (2020). Our results show a 27% Word Error Rate (WER) relative average improvement while requiring 45% less data than established baselines. We also adapt the standard WER to create an Uncertainty WER (U-WER) metric to track model adaptation to African accents.
The impact of our approach is substantial. It develops more robust, generalizable, and cost-efficient African-accented English ASR models, reducing dependency on large labeled datasets and enabling deployment in various real-world scenarios. Our results demonstrate improved generalization for out-of-distribution (OOD) cases, particularly for accents with limited resources, addressing specific challenges in African-accented automatic speech recognition (ASR). Additionally, by focusing on equitable representation in ASR training, our methodology promotes fairness in AI, ensuring technology serves users across diverse linguistic backgrounds without bias Selbst et al. (2019); Mitchell et al. (2019); Mehrabi et al. (2021). Our contributions are listed as follows:
  • we combine DBAL, AAL, CSA, and EU to propose a novel way to adapt several high-performing pretrained speech models to build efficient African-accented English ASR models,
  • we evaluate our approach across several speech domains (clinical, general, both), and African-accented speech datasets AfriSpeech-200 Olatunji et al. (2023b), SautiDB Afonja et al. (2021b), MedicalSpeech, and CommonVoices English Accented Dataset Ardila et al. (2019), while providing domain and accent-specific analyses,
  • we define a new and simple metric called U-WER that allows us to measure and track how the variance of the model, across hard accents, changes over the adaptation process,
  • we show that our approach improves the relative average WER performance by 27% while significantly reducing the required amount of labeled data (by $\sim$45%),
  • we show, based on additional AL experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
2 Background and Related Works
2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data Koenecke et al. (2020); Das et al. (2021). This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood Koenecke et al. (2020); Koenecke (2021); Chiu et al. (2018); Mengesha et al. (2021). Enhancing Automatic Speech Recognition (ASR) for African languages is crucial for achieving equitable voice recognition, particularly in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to the unique characteristics of these languages.
2.2 Active Learning
AL aims to reduce the number of labeled training examples by automatically processing unlabeled examples and selecting the most informative ones, considering a given cost function, for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization Settles (2009); Gal et al. (2017). Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset $\mathcal{D}_{\mathrm{pool}}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathrm{pool}}}$, a labeled training set $\mathcal{D}_{\mathrm{train}}=\{\mathbf{x}_{i},\mathrm{y}_{i}\}_{i=1}^{n_{\mathrm{train}}}$, and a predictive model with likelihood $p_{w}(y|x)$ parameterized by $w\sim p(W|\mathcal{D}_{\mathrm{train}})$ ($W$ are the parameters of the model). The setup assumes the presence of an oracle to provide predictions $\mathrm{y}$ for all $x_{i}\in\mathcal{D}_{\mathrm{pool}}$. After training, a batch of data $\{\mathbf{x}^{*}_{i}\}_{i=1}^{b}$ is selected from $\mathcal{D}_{\mathrm{pool}}$ based on its EU.
In Hakkani-Tür et al. (2002), AL was applied to a toy dataset of How May I Help You recordings. Confidence scores were estimated for each word and used to compute the overall confidence score for the audio sample. This approach achieved competitive results using 27% less data compared to the baseline. In Riccardi and Hakkani-Tur (2005), the authors estimated confidence scores for each utterance using an online algorithm with the lattice output of a speech recognizer. The utterance scores were filtered through an informativeness function to select an optimal subset of training samples, reducing the labeled data needed for a given WER by over 60%. Nallasamy et al. (2012) experimented with AL for accent adaptation in speech recognition. They adapted a source recognizer to the target accent by selecting a small, matched subset of utterances from a large, untranscribed, multi-accented corpus for human transcription. They employed a cross-entropy-based relevance measure in conjunction with uncertainty-based sampling. However, their experiments on Arabic and English accents showed worse performance compared to baselines while using more hours of recordings.
Figure 1: Our adaptation pipeline involves several phases. Initially, the dataset is split into a training set ($D1=\mathcal{D}_{\mathrm{train}}^{*}$, 30%) and a pool dataset ($D2=\mathcal{D}_{\mathrm{pool}}$, 70%). In the iterative process between phases 2 and 3, $D1$ is used to finetune a pretrained model. The top-k samples are selected using defined strategies and added to $D1$ for the next round. For more details on the uncertainty selection strategy, see section 3.2.
3 Datasets and Methodology
3.1 Datasets
We used the AfriSpeech-200 dataset Olatunji et al. (2023b), a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset comprises over 120 African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing the diversity of African regional languages. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US (see Table 1).
To demonstrate the dataset-agnostic nature of our approach, we also explored three additional datasets: (1) SautiDB Afonja et al. (2021a), Nigerian accent recordings with 919 audio samples at a 48kHz sampling rate, totaling 59 minutes; (2) MedicalSpeech [^1] https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent, containing 6,661 audio utterances of common medical symptoms, totaling 8 hours; and (3) CommonVoices English Accented Dataset, a subset of English Common Voice (version 10) Ardila et al. (2019), excluding western accents to focus on low-resource settings.
Table 1: AfriSpeech-200 Dataset statistics
AfriSpeech Dataset Statistics
Total duration 200.91 hrs
Total clips 67,577
Unique Speakers 2,463
Average Audio duration 10.7 seconds
Speaker Gender Ratios - # Clip %
Female 57.11%
Male 42.41%
Other/Unknown 0.48%
Speaker Age Groups - # Clips
<18yrs 1,264 (1.88%)
19-25 36,728 (54.58%)
26-40 18,366 (27.29%)
41-55 10,374 (15.42%)
>56yrs 563 (0.84%)
Clip Domain - # Clips
Clinical 41,765 (61.80%)
General 25,812 (38.20%)
Table 2: Dataset splits showing speakers, number of clips, and speech duration in Train/Dev/Test splits.
AfriSpeech-200 Dataset Splits
Item Train ($\mathcal{D}_{\mathrm{train}}^{*}$) Dev Test AL Top-k
# Speakers 1466 247 750
# Hours 173.4 8.74 18.77
# Accents 71 45 108
Avg secs/speaker 425.81 127.32 90.08
clips/speaker 39.56 13.08 8.46
speakers/accent 20.65 5.49 6.94
secs/accent 8791.96 698.82 625.55
# general domain 21682 (*6504) 1407 2723 2000
# clinical domain 36318 (*10895) 1824 3623 3500
# both domain 58000 (*17400) 3221 6346 6500
3.2 Methodology
Algorithm 1 Selection of the best-generated transcript in Active Learning for an input Sample $x$
1:we generate the predictions $\hat{y}_{1},..,\hat{y}_{T}$ corresponding to each stochastic forward pass ($T$=10 in our experiments)
2:we define a list variable called wer_list and a dictionary variable called wer_target_dict, respectively tracking all pairwise WERs and the average pairwise WER of each target prediction
3:for $\forall$ i,j $\in\{1,...,T\}$ do
4:     $\rightarrow$ $\hat{y}_{i}$ is set as target transcription
5:     $\rightarrow$ target_wer = list()
[…]
7:         $w=WER(\hat{y}_{j},\hat{y}_{i})$
8:         wer_list.append($w$)
9:         target_wer.append($w$)
10:     end for
11:     $wer_{\hat{y}_{i}}$ = mean(target_wer)
12:     wer_target_dict[$\hat{y}_{i}$] $\leftarrow$ $wer_{\hat{y}_{i}}$
13:end for
14:$\hat{y}_{best}$ = $\hat{y}_{i}$, such that wer_target_dict[$\hat{y}_{i}$] = min(wer_target_dict.values())
15:return ($p_{best}$, std(wer_list))
In our approach, to compute EU for a given input $x\in\mathcal{D}_{\mathrm{pool}}$, we perform MC-Dropout to obtain multiple stochastic forward passes through a finetuned ASR model $g$ with likelihood $p_{w\sim p(\mathbf{W}\mid\mathcal{D}_{\mathrm{train}}^{*})}(y|x)$ where $\mathbf{W}$ is the weights of $g$. Let $f$ be a function that computes the WER between the predicted and the target transcripts. Let $T$ be the number of stochastic forward passes. For each pass $t$, we apply dropout, obtain the output transcript, and compute the WER:
$f_{t}=f(y,\hat{y}_{t});\hat{y}_{t}=g(\mathbf{W},\tilde{x}_{t});\tilde{x}_{t}=x\cdot\mathbf{M}_{t}$
where $\mathbf{M}_{t}$ is a binary mask matrix sampled independently for each pass. EU($x|g,T$) can then be estimated from the $T$ stochastic forward passes as follows:
[…]
The use of MC-Dropout requires models to have dropout components during training. This exclusion applies to some models, such as Whisper Radford et al. (2022), which we still fine-tuned and evaluated as a baseline. We utilize four state-of-the-art pre-trained models: Wav2Vec2-XLSR-53, HuBERT-Large, WavLM-Large, and NVIDIA Conformer-CTC Large (en-US), referred to as Wav2Vec, HuBERT, WavLM, and Nemo, respectively.
3.2.1 Uncertainty WER
To handle diverse accents, we aim to reduce the EU of the models across hard accents after each adaptation round. We define a metric called U-WER to track this. To compute U-WER($a$) where $a$ is a hard accent, we condition EU on $a$:
[…]
where $x_{{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}}$ is the audio sample with accent a and
$f_{t,{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}}=f(y_{{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}},\hat{y}_{t,{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}});\hat{y}_{t,{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}}=g(\mathbf{W},\tilde{x}_{t,{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}});\tilde{x}_{t,{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}}=x_{{\color[rgb]{0.8359375,0.3125,0.078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8359375,0.3125,0.078125}a}}\cdot\mathbf{M}_{t}$
[…]
Algorithm 2 Adaptation Round using Epistemic Uncertainty-based Selection
1:Pretrained Model $\mathcal{M}$, Training Dataset $\mathcal{D}_{\mathrm{train}}^{*}$, Validation Dataset $\mathcal{D}_{Val}$, and Pool Dataset $\mathcal{D}_{\mathrm{pool}}$
2:$\mathcal{N}\leftarrow 3$$\triangleright$ Number of Adaptation Rounds
3:$T\leftarrow 10$$\triangleright$ Number of Stochastic Forward Passes
4:for $k\leftarrow 1$ to $\mathcal{N}$ do
5:     $g\leftarrow$ Finetune $\mathcal{M}$ on $\mathcal{D}_{\mathrm{train}}^{*}$ using $\mathcal{D}_{Val}$
6:     $\mathcal{EUL}\leftarrow\{\}$triangleright List of Uncertainty Scores
7:     for $x$ in $\mathcal{D}_{\mathrm{pool}}$ do $\triangleright$ x is an audio sample
8:         EUx $\leftarrow$ EU($x|g,T$) $\triangleright$ Epistemic Uncertainty of $x$
9:         $\mathcal{EUL}\leftarrow\mathcal{EUL}\cup\{(x,\text{EU}_{x})\}$
10:     end for
11:     $topk\leftarrow\{x_{1},...,x_{k}\}$triangleright Samples with highest $\mathcal{EU}$
12:     $\mathcal{D}_{\mathrm{train}}^{*}\leftarrow\mathcal{D}_{\mathrm{train}}^{*}\cup topk$
13:     $\mathcal{D}_{\mathrm{pool}}\leftarrow\mathcal{D}_{\mathrm{pool}}\setminus topk$
14:end for
3.3 Experimental Design
To work within our framework, we define the following selection strategies:
  • random: Randomly selects audio samples from $\mathcal{D}_{\mathrm{pool}}$.
  • EU-Most: Selects the most uncertain audio samples from $\mathcal{D}_{\mathrm{pool}}$ to add to $\mathcal{D}_{\mathrm{train}}$.
  • AL-EU-Most: Combines AL with the EU-Most strategy to finetune the pretrained model.
We also define standard fine-tuning (SFT) as baseline using all available data for finetuning. In SFT, $\mathcal{D}_{\mathrm{pool}}$ is empty. While running the defined strategies in our framework, we impose data constraints, not exceeding 60-65% of the initial dataset after all adaptation rounds. $\mathcal{D}_{\mathrm{train}}^{*}$ is 30% of $\mathcal{D}_{\mathrm{train}}$, and $\mathcal{D}_{\mathrm{pool}}$ is 70% of $\mathcal{D}_{\mathrm{train}}$. This simulates realistic scenarios where not all data might be available, testing the approach’s robustness and efficiency under constraints. The number of samples in $\mathcal{D}_{\mathrm{train}}$ and $\mathcal{D}_{\mathrm{pool}}$ is based on available training examples for each domain (see Tables 2, 4, and Appendix A.1).
Our EU-based pipeline is shown in Figure 1 and Algorithm 2. In each adaptation round, we use a finetuned model and a selection strategy to choose samples from $\mathcal{D}_{\mathrm{pool}}$ to add to $\mathcal{D}_{\mathrm{train}}^{*}$. During AL experiments, we consider samples from $\mathcal{D}_{\mathrm{pool}}$ as unlabeled: (1) using MC-Dropout, we obtain $n=10$ different input representations per audio sample to get $n$ different transcripts; (2) we then learn to select the best-generated transcription as the target transcription according to Algorithm 1.
Our experiments aim to answer the following research questions:
  1. 1.
    how does the pretrained ASR model adapt to a set of African accents across adaptation rounds and domains?
  2. 2.
    which selection strategy (EU-most or random) works better, and for which domain(s)?
  3. 3.
    which domain(s) help the model perform better, and how does the model perform (in terms of uncertainty) across the domain(s)?
  4. 4.
    what is the impact of EU-based selection on the model’s efficiency in low-resource data scenarios?
  5. 5.
    is uncertainty-based selection, model, and dataset agnostic?
U-WER will answer question 4. To answer question 5, we evaluated our approach with three additional pretrained models (Nemo, WavLM, and Hubert) and across three external datasets (SautiDB, CommonVoices English Accented Dataset, and MedicalSpeech). For consistency and better visualization, we considered the top-10 (in terms of frequency) accents across three adaptation rounds and both selection strategies to answer questions 1-4. For very low-resource settings, we considered the five accents with the least recording hours.
For our experiments, we utilized six RTX 8000 GPUs and four A100 GPUs. Training and evaluation were conducted over a period of one month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16 kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
Table 3: We utilized Wav2Vec to conduct initial experiments across various domains and strategies, aiming to identify the optimal selection strategy. Models marked with ** are used to demonstrate that our algorithm is model agnostic, utilizing the EU-Most selection strategy, which has been proven to be the most effective. Our AL experiments also use this strategy. Wav2Vec, using the random strategy, scored 0.1111, 0.3571, and 0.1666 for the general, clinical, and both domains, respectively. We omit random results to enhance readability.
Model General Clinical Both
Baseline EU-Most AL-EU-Most Baseline EU-Most AL-EU-Most Baseline EU-Most AL-EU-Most
Wav2vec 0.2360 Olatunji et al. (2023b) 0.1011 0.1059 0.3080 Olatunji et al. (2023b) 0.2457 0.2545 0.2950 Olatunji et al. (2023b) 0.1266 0.1309
**Hubert 0.1743 0.1901 0.1887 0.2907 0.2594 0.2709 0.2365 0.2453 0.2586
**WavLM 0.1635 0.1576 0.1764 0.3076 0.2313 0.2537 0.2047 0.1897 0.1976
*Nemo 0.2824 0.1765 0.1815 0.2600 0.2492 0.2526 0.3765 0.2576 0.2610
Average Performance 0.2141 0.1563 0.1631 0.2916 0.2464 0.2579 0.2782 0.2043 0.2120
Whisper-Medium 0.2806 - - 0.3443 - - 0.3116 - -
Figure 2: WER Performance on Accents from General Domain
Figure 3: WER Performance on Accents from Clinical Domain
Figure 4: WER Performance on Accents from Clinical+General (Both) Domain
Table 4: WER Evaluation Results on External Datasets, with $\alpha\in[0.60,0.65]$ as described in Section 3.1 and on Figure 1. We observe an improvement in WER using our approach across all datasets, indicating that our algorithm is dataset-agnostic.
Dataset Split and Size for our approach Finetuning Epochs Baseline EU-Most
$\mathcal{D}_{\mathrm{train}}^{*}$ $\mathcal{D}_{\mathrm{pool}}$ Top-k Test ($\mathcal{D}_{\mathrm{train}}$) ($\mathcal{D}_{\mathrm{train}}^{*}$ + $\alpha\mathcal{D}_{\mathrm{pool}}$)
SautiDB Afonja et al. (2021a) 234 547 92 138 50 0.50 0.12
MedicalSpeech 1598 3730 1333 622 5 0.30 0.28
CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) 26614 62100 10350 232 5 0.50 0.22
Average 0.43 0.20
4 Results and Discussion
To assess the performance improvement for each domain, we compute the relative average improvement
$\text{RIA}_{wer,d}=\left(\frac{b_{wer}^{d}-s_{wer}^{d}}{b_{wer}^{d}}\right)\times 100\%$
where $b_{wer}^{d}$ and $s_{wer}^{d}$ are the average WER respectively of the baseline, and the best selection strategy, in a domain $d\in\{general,clinical,both\}$. A higher percentage reflects a higher improvement in our approach.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (27.00%), clinical (15.51%), and both (26.56%). Our approach also surpasses Whisper-Medium (Olatunji et al. (2023b); Radford et al. (2023)), demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model’s exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between the general and clinical domains are noted, likely due to the complexity of the clinical sample. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, facilitate model learning and improve performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the limited resource availability typical of many African languages and dialects.
To demonstrate the agnostic aspect of our approach, we evaluated it using three additional pre-trained models (Hubert, WavLM, and Nemo) and three datasets containing accented speech in general and clinical domains, employing only the EU-Most selection strategy. The results, shown in Tables 3 and 4, indicate that our uncertainty-based adaptation approach consistently outperforms baselines. This confirms that our approach applies to any model architecture and dataset, allowing us to answer question 5 positively.
5 Conclusion
We combined several AL paradigms, the CSA, and the EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced the U-WER metric to track model adaptation to intricate accents. Our experiments demonstrated a remarkable 27% WER ratio improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This reflects the efficiency and potential of our approach to lower the barriers to ASR technologies in underserved regions significantly. Our method enhances model robustness and generalization across various domains, datasets, and accents, which are crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
6 Limitations
In discussing trade-offs (Section 4), we noted that while our approach enhances performance, particularly with linguistically rich accents, a stopping criterion is essential for complex domains like the clinical one to balance adaptation rounds with the pool size. With better resources, we would consider implementing Deep Ensembles (Lakshminarayanan et al. (2017)) as an alternative to our current MC-Dropout method for estimating epistemic uncertainty and leveraging other acquisition functions (such as BALD, BatchBALD) highlighted in this work.
7 Acknowledgments
The authors are very grateful to Prof. Ines Arous for her help in reviewing the manuscript and suggesting very important changes to improve the quality of the paper. The author acknowledges the support of the Mila Quebec AI Institute for computing resources.
The authors acknowledge Intron Health for providing the Afrispeech-200 dataset. The authors are grateful to Atnafu Lambebo Tonja, Chris Chinenye Emezue, Tobi Olatunji, Naome A Etori, Salomey Osei, Tosin Adewumi, and Sahib Singh for their help in the early stage of this project.
Appendix A Appendices
A.1 Hyper-parameters
Table 5 shows the hyper-parameter settings used in this study. The top-k value in the table is changed according to the domain used in each of the experiments. For example, when conducting experiments in the general domain, we set the value of top-k to 2k.
Hyper-parameters Values
attention dropout 0.1
hidden dropout 0.1
layer drop 0.1
train batch size 16
val batch size 8
number of epochs 5
learning rate 3e-4
maximum audio length 260000
maximum label length 260
minimum transcript length 10
top_k 2000, 3500, 6500
domains general, clinical, all
active learning rounds 3
sampling mode EU-Most, random
MC-Dropout round 10
Table 5: Hyper-parameters summary
A.2 Country Statistics
Table 6 shows the countries’ statistics across the AfriSpeech-200 dataset.
Country Clips Speakers Duration (seconds) Duration (hrs)
Nigeria 45875 1979 512646.88 142.40
Kenya 8304 137 75195.43 20.89
South Africa 7870 223 81688.11 22.69
Ghana 2018 37 18581.13 5.16
Botswana 1391 38 14249.01 3.96
Uganda 1092 26 10420.42 2.89
Rwanda 469 9 5300.99 1.47
United States of America 219 5 1900.98 0.53
Turkey 66 1 664.01 0.18
Zimbabwe 63 3 635.11 0.18
Malawi 60 1 554.61 0.15
Tanzania 51 2 645.51 0.18
Lesotho 7 1 78.40 0.02
Table 6: Countries Statistics across the dataset
A.3 Dataset Accents Stats
Tables 7 and 8 provide a list of AfriSpeech accents along with the number of unique speakers, countries where speakers for each accent are located, duration in seconds for each accent, and their presence in the train, dev, and test splits.
Accent Clips Speakers Duration(s) Countries Splits
yoruba 15407 683 161587.55 US,NG train,test,dev
igbo 8677 374 93035.79 US,NG,ZA train,test,dev
swahili 6320 119 55932.82 KE,TZ,ZA,UG train,test,dev
hausa 5765 248 70878.67 NG train,test,dev
ijaw 2499 105 33178.9 NG train,test,dev
afrikaans 2048 33 20586.49 ZA train,test,dev
idoma 1877 72 20463.6 NG train,test,dev
zulu 1794 52 18216.97 ZA,TR,LS dev,train,test
setswana 1588 39 16553.22 BW,ZA dev,test,train
twi 1566 22 14340.12 GH test,train,dev
isizulu 1048 48 10376.09 ZA test,train,dev
igala 919 31 9854.72 NG train,test
izon 838 47 9602.53 NG train,dev,test
kiswahili 827 6 8988.26 KE train,test
ebira 757 42 7752.94 NG train,test,dev
luganda 722 22 6768.19 UG,BW,KE test,dev,train
urhobo 646 32 6685.12 NG train,dev,test
nembe 578 16 6644.72 NG train,test,dev
ibibio 570 39 6489.29 NG train,test,dev
pidgin 514 20 5871.57 NG test,train,dev
luhya 508 4 4497.02 KE train,test
kinyarwanda 469 9 5300.99 RW train,test,dev
xhosa 392 12 4604.84 ZA train,dev,test
tswana 387 18 4148.58 ZA,BW train,test,dev
esan 380 13 4162.63 NG train,test,dev
alago 363 8 3902.09 NG train,test
tshivenda 353 5 3264.77 ZA test,train
fulani 312 18 5084.32 NG test,train
isoko 298 16 4236.88 NG train,test,dev
akan (fante) 295 9 2848.54 GH train,dev,test
ikwere 293 14 3480.43 NG test,train,dev
sepedi 275 10 2751.68 ZA dev,test,train
efik 269 11 2559.32 NG test,train,dev
edo 237 12 1842.32 NG train,test,dev
luo 234 4 2052.25 UG,KE test,train,dev
kikuyu 229 4 1949.62 KE train,test,dev
bekwarra 218 3 2000.46 NG train,test
isixhosa 210 9 2100.28 ZA train,dev,test
hausa/fulani 202 3 2213.53 NG test,train
epie 202 6 2320.21 NG train,test
isindebele 198 2 1759.49 ZA train,test
venda and xitsonga 188 2 2603.75 ZA train,test
sotho 182 4 2082.21 ZA dev,test,train
akan 157 6 1392.47 GH test,train
nupe 156 9 1608.24 NG dev,train,test
anaang 153 8 1532.56 NG test,dev
english 151 11 2445.98 NG dev,test
afemai 142 2 1877.04 NG train,test
shona 138 8 1419.98 ZA,ZW test,train,dev
eggon 137 5 1833.77 NG test
luganda and kiswahili 134 1 1356.93 UG train
ukwuani 133 7 1269.02 NG test
sesotho 132 10 1397.16 ZA train,dev,test
benin 124 4 1457.48 NG train,test
kagoma 123 1 1781.04 NG train
nasarawa eggon 120 1 1039.99 NG train
tiv 120 14 1084.52 NG train,test,dev
south african english 119 2 1643.82 ZA train,test
borana 112 1 1090.71 KE train
Table 7: Dataset Accent Stats, Part I
Accent Clips Speakers Duration(s) Countries Splits
swahili ,luganda ,arabic 109 1 929.46 UG train
ogoni 109 4 1629.7 NG train,test
mada 109 2 1786.26 NG test
bette 106 4 930.16 NG train,test
berom 105 4 1272.99 NG dev,test
bini 104 4 1499.75 NG test
ngas 102 3 1234.16 NG train,test
etsako 101 4 1074.53 NG train,test
okrika 100 3 1887.47 NG train,test
venda 99 2 938.14 ZA train,test
siswati 96 5 1367.45 ZA dev,train,test
damara 92 1 674.43 NG train
yoruba, hausa 89 5 928.98 NG test
southern sotho 89 1 889.73 ZA train
kanuri 86 7 1936.78 NG test,dev
itsekiri 82 3 778.47 NG test,dev
ekpeye 80 2 922.88 NG test
mwaghavul 78 2 738.02 NG test
bajju 72 2 758.16 NG test
luo, swahili 71 1 616.57 KE train
dholuo 70 1 669.07 KE train
ekene 68 1 839.31 NG test
jaba 65 2 540.66 NG test
ika 65 4 576.56 NG test,dev
angas 65 1 589.99 NG test
ateso 63 1 624.28 UG train
brass 62 2 900.04 NG test
ikulu 61 1 313.2 NG test
eleme 60 2 1207.92 NG test
chichewa 60 1 554.61 MW train
oklo 58 1 871.37 NG test
meru 58 2 865.07 KE train,test
agatu 55 1 369.11 NG test
okirika 54 1 792.65 NG test
igarra 54 1 562.12 NG test
ijaw(nembe) 54 2 537.56 NG test
khana 51 2 497.42 NG test
ogbia 51 4 461.15 NG test,dev
gbagyi 51 4 693.43 NG test
portuguese 50 1 525.02 ZA train
delta 49 2 425.76 NG test
bassa 49 1 646.13 NG test
etche 49 1 637.48 NG test
kubi 46 1 495.21 NG test
jukun 44 2 362.12 NG test
igbo and yoruba 43 2 466.98 NG test
urobo 43 3 573.14 NG test
kalabari 42 5 305.49 NG test
ibani 42 1 322.34 NG test
obolo 37 1 204.79 NG test
idah 34 1 533.5 NG test
bassa-nge/nupe 31 3 267.42 NG test,dev
yala mbembe 29 1 237.27 NG test
eket 28 1 238.85 NG test
afo 26 1 171.15 NG test
ebiobo 25 1 226.27 NG test
nyandang 25 1 230.41 NG test
ishan 23 1 194.12 NG test
bagi 20 1 284.54 NG test
estako 20 1 480.78 NG test
gerawa 13 1 342.15 NG test
Table 8: Dataset Accent Stats, Part II
A.4 Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with random and EU-Most selection strategies.
Figure 5: Most common accents distribution across the general domain with EU-Most sampling strategy.
Figure 6: Most common accents distribution across the general domain with random selection strategy.
A.5 Ascending and Descending Accents
Figure 7 shows ascending and descending accents across the Top 2k most uncertain samples.
Figure 7: Ascending and descending accents across Top-2K most uncertain samples.